Find the equation of the line tangent to the circle x2 + y2 = 36 at point (11,5). Use the general equation of the line for your final answer.
Answer by Boreal(15235)

The problem here is that the point (11, 5) is NOT on the circle but (sqrt (11), 5) is
x^2+y^2=36
implicit differentiation
2x+2y dy/dx=0
2y dy/dx=-2x
dy/dx=-x/y, and that is the slope of the tangent line to the circle.
calling it m, m=-sqrt(11)/5
point slope formula y-y1=m(x-x1), m slope, (x1, y1) point
y-5=(-sqrt(11)/5)(x-sqrt (11))
y-5=-(sqrt(11)/5)x+(11/5)
y=-(sqrt(11)/5)x+7.2